OpenStudy (anonymous):

I have an equation in the format of y=ax^2 y=38 x=360 I have it set up as follows: 38=a(180)^2 I am stuck at 38/32400=19/16200 is this the lowest? Then i get 19/16200x^2 Is this correct?

OpenStudy (anonymous):

y=ax^2 38=a times 360^2

OpenStudy (anonymous):

can you figure the rest out?

OpenStudy (anonymous):

38= a times 129600

OpenStudy (anonymous):

if you want to find a, a= 38/129600

OpenStudy (anonymous):

hasteeerfani00 I must have multiplied wrong

OpenStudy (anonymous):

Would it not be 180^2 which is 32400?

OpenStudy (anonymous):

180^2 means 180 squared. so it would be 32400

OpenStudy (anonymous):

the question says x=360

OpenStudy (anonymous):

yes and 38/32400 gives me a decimal I think it has to be in fraction form so I can put it like this 1/675x^2 (different example)

OpenStudy (anonymous):

i can reduce it to 19/16200 but thats it. Is this the answer? 19/16200x^2

OpenStudy (anonymous):

well i beileve that you would have to have plugged x in

OpenStudy (anonymous):

what is this for? geometry, algebra etc

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

one or two

OpenStudy (anonymous):

I had another problem and i did nto get decimals. it was 48=a(180)^2 Then I got 48/32400=1/675 Therefore I got y=1/675x^2 But thid problem is giving me a decimal

OpenStudy (anonymous):

college algebra